OpenStudy (anonymous):

find the area of a regular hexagon with an apothem 10.4yards and a side 12 yards long. Round your answer to the nearest tenth

OpenStudy (michele_laino):

we have to apply this formula: \[\Large area = \frac{{perimeter \times apothem}}{2}\]

OpenStudy (michele_laino):

furthermore the perimeter is: \[\Large perimeter = side \times 6=12 \times 6=...?\]

OpenStudy (michele_laino):

the side is 12 yard, as I can read from the text of your problem

OpenStudy (anonymous):

or)) you are given the size of one triangle base 12 height 10.4 so area =1/2 (12)(10.4) =62.4 now you have 6 same triangles so 6(62.4)=374.4